OpenStudy (anonymous):

What is the simplest form for a to the power of 9 b to the power of 5 over a to the power of 3 b to the power of 15

OpenStudy (campbell_st):

can you draw the powers

OpenStudy (anonymous):

Is that what you're trying to say? \[(A^{9} B^{5} ) \div (A^{3} B^{15})\]

OpenStudy (anonymous):

@Wired yes! :)

OpenStudy (anonymous):

@Wired

OpenStudy (campbell_st):

when dividing subtract the powers for the same base \[\frac{x^a}{x^b} = x^{a -b}\] so you have \[a^{9-3}\times b^{5 - 15}\] negative powers indicate fractions \[x^{-b} = \frac{1}{x^b}\] hope this helps

OpenStudy (anonymous):

@Wired helppleaseeee

OpenStudy (anonymous):

campbell already gave you the proper equations.